OpenStudy (anonymous):

solve: log (x+1) + log (x+4)=1

OpenStudy (anonymous):

First combine the two logs via the multiplication rule.

OpenStudy (anonymous):

log((x+1)(x+4)) = 1

OpenStudy (anonymous):

Then convert to exponent form. (the base is 10 on this one)

OpenStudy (anonymous):

so i did X^2+5x+3=0 and stuck after that

OpenStudy (anonymous):

+3???

OpenStudy (anonymous):

should be -6

OpenStudy (anonymous):

\[x ^{2}+5x-6\]

OpenStudy (anonymous):

The base is 10

OpenStudy (anonymous):

how did you get -6?

OpenStudy (anonymous):

so 10 = (x+1)(x+4)

OpenStudy (anonymous):

the solutions must be >-1

OpenStudy (anonymous):

That make sense?

OpenStudy (anonymous):

ohh ok! but what would i do with the 1?

OpenStudy (anonymous):

The 1 is your exponent man.

OpenStudy (anonymous):

\[10^{1}=(x+1)(x+4)\]

OpenStudy (anonymous):

Just simplify that 1 out of the equation...you know.

OpenStudy (anonymous):

i wrote it down. i got it. let me try solving it and ill get back to you

OpenStudy (anonymous):

Sounds good man.

OpenStudy (anonymous):

Don't forget to mark a Best Response and take care...I'll be around for little while longer, if you need something else just holler.

OpenStudy (anonymous):

you will find two solutions...one of them is not match !

OpenStudy (anonymous):

so i got the solutions x=-6 and x=1 but we can't use negatives in log equations right so the answer would be 1?

OpenStudy (anonymous):

yeahh ...true...x must be x>-1 ! so the only solution is 1

OpenStudy (anonymous):

Good job man

OpenStudy (anonymous):

thanks for all your help!

OpenStudy (anonymous):

Anytime. Take care!

OpenStudy (anonymous):

you're welcome :) !

OpenStudy (anonymous):

can you guys help me with one more problem please? \[3e ^{3x-1}-2=13\]

OpenStudy (anonymous):

You got lucky man. 1) Add 2 to both sides.

OpenStudy (anonymous):

so i know my first step is to add 2

OpenStudy (anonymous):

2) Divide by 3

OpenStudy (anonymous):

ok did that

OpenStudy (anonymous):

take the natural log (ln) of both sides.

OpenStudy (anonymous):

...in order to bring down the exponent on the left.

OpenStudy (anonymous):

and eliminates the e

OpenStudy (anonymous):

This give: 3x - 1 = ln(5)

OpenStudy (anonymous):

ok so would it just be 3x-1=ln5

OpenStudy (anonymous):

Add 1. Divide by 3.

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

ok cool! i know where i messed up. Thanks again for your help :)

OpenStudy (anonymous):

No prob. Later